OpenStudy (anonymous):

Need help with system: -4x-2y=14 -10x+7y=-25

OpenStudy (anonymous):

what are you trying to do? substitute?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

-4x-2y=14 Since -4x does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4x to both sides. -2y=4x+14 Divide each term in the equation by -2. -(2y)/(-2)=(4x)/(-2)+(14)/(-2) Simplify the left-hand side of the equation by canceling the common factors. y=(4x)/(-2)+(14)/(-2) Simplify the right-hand side of the equation by simplifying each term. y=-2x-7

OpenStudy (anonymous):

-10x+7y=-25 Since -10x does not contain the variable to solve for, move it to the right-hand side of the equation by adding 10x to both sides. 7y=10x-25 Divide each term in the equation by 7. (7y)/(7)=(10x)/(7)-(25)/(7) Simplify the left-hand side of the equation by canceling the common factors. y=(10x)/(7)-(25)/(7) Simplify the right-hand side of the equation by simplifying each term. y=(5(2x-5))/(7)